OpenStudy (anonymous):

In “The Leap,” the narrator’s mother deals with the past in an unusual way. She.

sammixboo (sammixboo):

\(\mathbb{ \Huge Welcome~to~\color{aqua}{Open}\color{olive}{Study!}}\) \(\Large\mathbb{Do~ you~ still ~need~\color{aqua}{help?}}\) \(\Large \mathbb{I~see~you~didn't~post~your~full~question!}\)